# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 4.2 Extending Theorem 1 to amenable groups
Definition 4.6. For an amenable semigroup $G$ with Følner sequence $\mathbf{F}$ and $S\subseteq G$ let
$\Delta_{1}(S)=\{g\in G:S\cap(g^{-1}S)\neq\emptyset\}$ (4.10)
$\Delta_{2}(\mathbf{F},S)=\{g\in G:\overline{d}_{\mathbf{F}}(S\cap(g^{-1}S))>0\}$ (4.11)
Remark 4.7.
  1. i)
    When $G$ is a group, the occurrences of $S\cap(g^{-1}S)$ in equations (4.10) and (4.11) can be replaced by $S\cap(gS)$ without changing the meaning of the Definition 4.2. In other words, when $G$ is a group, both $\Delta_{1}(S)$ and $\Delta_{2}(\mathbf{F},S)$ are symmetric, meaning that $g\in\Delta_{1}(S)\iff g^{-1}\in\Delta_{1}(S)$, and $g\in\Delta_{2}(\mathbf{F},S)\iff g^{-1}\in\Delta_{2}(\mathbf{F},S)$.
  2. ii)
@@
Theorem 4.8 (Theorem 1 for countably infinite amenable groups). Let $G$ be a countably infinite amenable group with (left) Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ where $\mathbf{G}$ is (left) tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$ where
@@
so $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$. Let $D$ denote the set $\bigcap_{i=1}^{k}{\Delta_{2}(\mathbf{F}_{i}},S_{i})$. In order to show that $\Delta_{1}(S)\subseteq D$, take $s,t\in S$ and note that
$0<\mu(T_{s}^{-1}B\cap T_{t}^{-1}B)$ $=\mu(B\cap T_{ts^{-1}}^{-1}B)$
$=\prod_{i=1}^{k}\mu_{i}(B_{i}\cap T_{i,ts^{-1}}^{-1}B_{i})\leq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(ts^{-1})^{-1}S_{i}).$
@@
It follows that $ts^{-1}\in\Delta_{2}(\mathbf{F},S_{i})$ for all $i\in\{1,\ldots,k\}$, meaning that $ts^{-1}\in D$. Thus $\Delta_{1}(S)\subseteq D$, as desired. Lastly, there are $m_{1},\ldots,m_{\ell}\in G$ where
$\ell\leq\prod_{i=1}^{k}1/\overline{d}_{\mathbf{F}_{i}}(S_{i})$
@@
such that $G=m_{1}\Delta_{1}(S)\cup\cdots\cup m_{\ell}\Delta_{1}(S)=\Delta_{1}(S)m_{1}^{-1}\cup\cdots\cup\Delta_{1}(S)m_{\ell}^{-1}$ (this will be proven below in Theorem 4.2). Since $\Delta_{1}(S)\subset D$ it follows that $G=m_{1}D\cup\cdots\cup m_{\ell}D=Dm_{1}^{-1}\cup\cdots\cup Dm_{\ell}^{-1}$. ∎
@@
The set $S$ in Theorem 4.7 has the property that $\Delta_{1}(S)\subset\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$ which means that for all $g$ such that $S\cap g^{-1}S\neq\emptyset$ it is true that $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g^{-1}S_{i}))>0$ for $1\leq i\leq k$. In fact one can show $S$ satisfies the stronger property that for all $g_{1},\dots,g_{n}\in G$ such that $S\cap(g_{1}^{-1}S)\cap\cdots\cap(g_{n}^{-1}S)\neq\emptyset$ it is true that $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g_{1}^{-1}S_{i})\cap\cdots\cap(g_{n}^{-1}S_{i}))>0$ for $1\leq i\leq k$. This result is reflected in the following theorem
@@
Theorem 4.9. Let $G$ be a countably infinite amenable group with (left) Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ where $\mathbf{G}$ is (left) tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is a set $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and if $g_{1},\ldots,g_{n}\in G$ are such that $S\cap(g_{1}^{-1}S)\cap\cdots\cap(g_{n}^{-1}S)\neq\emptyset$, then $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g_{1}^{-1}S_{i})\cap\cdots\cap(g_{n}^{-1}S_{i}))>0$ for $1\leq i\leq k$.
@@
Utilizing the fact that every Følner sequence has a tempered Følner subsequence (see [13, Proposition 1.4]), it is straightforward to obtain the following variant of Theorem 4.2 with the slightly weaker assumption that the Følner sequence $\mathbf{G}$ may not be tempered.
@@
Theorem 4.10. Let $G$ be a countably infinite amenable group and let $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $G$. If $S_{1},\ldots,S_{k}\subseteq G$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there is $S\subseteq G$ such that $\overline{d}_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$ where
@@
To complete the logical circle we still owe the reader the proof of Theorem 4.2 which was used in the proof of Theorem 4.2. Before proceeding, we need to introduce some preliminary definitions and prove some necessary lemmas. Let $G$ be a semigroup and let $S\subseteq G$. The set $S$ is left syndetic if there are $c_{1},\ldots,c_{k}\in G$ such that $c_{1}^{-1}S\cup\cdots\cup c_{k}^{-1}S=G$, similarly $S$ is right syndetic if there are $d_{1},\dots,d_{r}\in G$ such that $Sd_{1}^{-1}\cup\dots\cup Sd_{r}^{-1}=G$. $S$ is right thick if for every finite $F\subseteq G$, there exists $g\in G$ for which $Fg\subseteq S$.
@@
Proof.
If $T=G$, then the result is clear. So suppose that $T\neq G$ and fix $t\in G\setminus T$. We inductively construct an injective sequence $s_{1},s_{2},...$ in $G$ with our desired property. Pick arbitrary $s_{1}\in G$. Now let $n\in\mathbb{N}$ be arbitrary and suppose that we have chosen pairwise distinct $s_{1},...,s_{n}\in G$ with $\{s_{i}^{-1}s_{j}:1\leq i<j\leq n\}\subseteq T$. Now $F=\left\{s_{1}^{-1},...,s_{n}^{-1},ts_{1}^{-1},...,ts_{n}^{-1}\right\}$ is finite, so take $s_{n+1}\in G$ such that $Fs_{n+1}\subseteq T$. Since $t\not\in T$, we have $s_{n+1}\neq s_{j}$ for any $1\leq j\leq n$. Furthermore one can now see that $\{s_{i}^{-1}s_{j}:1\leq i<j\leq n+1\}\subseteq T$, so by induction this gives us an infinite sequence $s_{1},s_{2},...$ with the desired properties ∎
@@
Proof. Let $F\subseteq G$ be finite and non-empty, say $F=\{g_{1},\ldots,g_{k}\}$. Then $g_{1}^{-1}S\cup\cdots\cup g_{k}^{-1}S\neq G$ by non-syndeticity of $S$, so pick $a\in G\setminus(g_{1}^{-1}S\cup\cdots\cup g_{k}^{-1}S)$. It follows that for all $g\in F$ we have $ga\notin S$, hence $Fa\subseteq G\setminus S=T$, which implies that $T$ is right thick as desired. ∎
Lemma 4.13. Let $G$ be a countably infinite amenable group and let $\mathbf{F}$ be a Følner sequence in $G$. Let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. If $(s_{n})_{n=1}^{\infty}$ is an injective sequence in $G$ then there are $i<j$ such that $s_{i}^{-1}s_{j}\in\Delta_{1}(S)$.
Proof. Let $(s_{n})_{n=1}^{\infty}$ be an injective sequence in $G$. By the pigeonhole principle, there must be some $i<j$ for which $\overline{d}_{\mathbf{F}}((s_{i}S)\cap(s_{j}S))>0$. It follows that there are $u,v\in S$ such that $s_{j}u=s_{i}v$, thus $(s_{i}^{-1}s_{j})u=v$, and therefore $u\in S\cap(s_{i}^{-1}s_{j})^{-1}S$. It follows that $S\cap(s_{i}^{-1}s_{j})^{-1}S\neq\emptyset$, therefore $s_{i}^{-1}s_{j}\in\Delta_{1}(S)$. ∎
We can now prove Theorem 4.2 completing hereby the proof of Theorem 4.2
Theorem 4.14. Let $G$ be a countably infinite amenable group, let $\mathbf{F}$ be a Følner sequence in $G$, and let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. Let $E=\Delta_{1}(S)$. Then there exist $b_{1},...,b_{\ell}\in G$ with $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$ for which $b_{1}E\cup\cdots\cup b_{\ell}E=Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}=G$ (that is, $E$ is both left and right syndetic with the same ”syndeticity constant” $\ell$).
Proof. By Lemma 4.2, for every injective sequence $(s_{n})_{n=1}^{\infty}$ in $G$ there are $i<j$ such that $s_{i}^{-1}s_{j}\in E$. Then by Lemma 4.2 it follows that $G\setminus E$ cannot be right thick, so $E$ is left syndetic by Lemma 4.2, so there are $c_{1},...,c_{k}\in G$ such that $c_{1}E\cup\cdots\cup c_{k}E=G$.
For any additional $g\in G$ with $g\notin\{c_{1},\ldots,c_{k}\}$ we have $g\in G=c_{1}E\cup\cdots\cup c_{k}E$, so $g\in c_{j}E$ for some $j\in\left\{1,...,k\right\}$. Hence $c_{j}^{-1}g\in E$, which means that $S\cap(c_{j}^{-1}g)^{-1}S\neq\emptyset$, so there are $u,v\in S$ such that $c_{j}^{-1}gu=v$. Thus $c_{j}\{u,v\}\cap g\{u,v\}\neq\emptyset$. It follows that there is a maximal subset $\left\{b_{1},...,b_{\ell}\right\}$ in $G$ such that, for all finite $F\subseteq S$, we have $b_{i}F\cap b_{j}F=\emptyset$ when $i\neq j$, and if $g\notin\{b_{1},\ldots,b_{\ell}\}$ then there is some finite $F\subseteq S$ such that $b_{i}F\cap gF\neq\emptyset$ for some $1\leq i\leq\ell$.
Let $\mathbf{F}=(F_{n})_{n=1}^{\infty}$ and for all $n\in\mathbb{N}$ put $S(n)=|A_{n}|$, where $A_{n}=S\cap F_{n}$. Now by assumption, for all $n\in\mathbb{N}$ the sets $b_{1}A_{n},...,b_{\ell}A_{n}$ are disjoint. Hence,
@@
By taking $n\to\infty$ in equation (4.14) and using the fact that F is a Følner sequence, it follows that $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$.
@@
Lastly, note that if $d\in E$, then $d^{-1}\in E$. For all $g\in G$ we have $g^{-1}\in b_{1}E\cup\cdots\cup b_{\ell}E$, hence $b_{i}^{-1}g^{-1}=d$ for some $d\in E$ and so $gb_{i}=d^{-1}\in E$, thus $g\in Eb_{i}^{-1}$. This shows that $G=Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}$. ∎
4.3 Extending Theorem 1 to Cancellative Amenable Semigroups
In this subsection we will obtain a variant of Theorem 4.2 for cancellative amenable semigroups, namely Theorem 4.3. We will do so by utilizing the classical fact that cancellative amenable semigroups can be embedded into groups. Let $G$ be a cancellative amenable semigroup. It is known that there is a group $\widetilde{G}$ and an embedding $\varphi\colon G\to\widetilde{G}$ such that $\widetilde{G}=\left\{\varphi(s)\varphi(t)^{-1}:s,t\in G\right\}$ (see [11, Theorem 1.23] and [14, Proposition 1.23]). Moreover, the group $\widetilde{G}$ and the emebedding $\varphi$ are unique (see [11, Theorem 1.25] for the precise formulation). With this in mind, we henceforth identify every element $g\in G$ with $\varphi(g)\in\widetilde{G}$ and call $\widetilde{G}$ the group of quotients of $G$.
@@ 4.3 Extending Theorem 1 to Cancellative Amenable Semigroups
Theorem 4.15 ([1, Theorem 2.12]). Let $G$ be a countably infinite amenable cancellative semigroup. Then any Følner sequence $(F_{N})_{N=1}^{\infty}$ in $G$ is a Følner sequence in its group of quotients $\widetilde{G}$.
Remark 4.16. Let $G$ be a countably infinite amenable cancellative semigroup and let $\mathbf{F}=(F_{N})_{N=1}^{\infty}$ be a Følner sequence in $G$, and hence in $\widetilde{G}$. Let $S^{\prime}\subset\widetilde{G}$ and let $S\vcentcolon=S^{\prime}\cap G$. We have
@@
Moreover, if one of $d_{\mathbf{F}}(S)$ or $d_{\mathbf{F}}(S^{\prime})$ exists then both exist and $d_{\mathbf{F}}(S)=d_{\mathbf{F}}(S^{\prime})$.

### v2

@@ 4.2 Extending Theorem 1 to amenable groups
Definition 4.7. For an amenable semigroup $G$ with Følner sequence $\mathbf{F}$ and $S\subseteq G$, let
$\Delta_{1}(S)=\{g\in G:S\cap(g^{-1}S)\neq\emptyset\},$ (4.10)
$\Delta_{2}(\mathbf{F},S)=\{g\in G:\overline{d}_{\mathbf{F}}(S\cap(g^{-1}S))>0\}.$ (4.11)
Remark 4.8. When $G$ is a group, the occurrences of $S\cap(g^{-1}S)$ in equations (4.10) and (4.11) can be replaced by $S\cap(gS)$ without changing the meaning of the Definition 4.2. In other words, when $G$ is a group, both $\Delta_{1}(S)$ and $\Delta_{2}(\mathbf{F},S)$ are symmetric, meaning that $g\in\Delta_{1}(S)\iff g^{-1}\in\Delta_{1}(S)$ and $g\in\Delta_{2}(\mathbf{F},S)\iff g^{-1}\in\Delta_{2}(\mathbf{F},S)$.
@@
Theorem 4.9 (Theorem 1 for countably infinite amenable groups). Let $G$ be a countably infinite amenable group with (left) Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$, where $\mathbf{G}$ is (left) tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$, where
@@
so $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$. Recall $D$ is defined to be the set $\bigcap_{i=1}^{k}{\Delta_{2}(\mathbf{F}_{i}},S_{i})$. In order to show that $\Delta_{1}(S)\subseteq D$, take $s,t\in S$ and note that
$0<\mu(T_{s}^{-1}B\cap T_{t}^{-1}B)$ $=\mu(B\cap T_{ts^{-1}}^{-1}B)=\prod_{i=1}^{k}\mu_{i}(B_{i}\cap T_{i,ts^{-1}}^{-1}B_{i})\leq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(ts^{-1})^{-1}S_{i}).$
@@
It follows that $ts^{-1}\in\Delta_{2}(\mathbf{F},S_{i})$ for all $i\in\{1,\ldots,k\}$, meaning that $ts^{-1}\in D$. Thus $\Delta_{1}(S)\subseteq D$, as desired. Lastly, there are $m_{1},\ldots,m_{\ell}\in G$, where
$\ell\leq\prod_{i=1}^{k}1/\overline{d}_{\mathbf{F}_{i}}(S_{i}),$
@@
such that $G=m_{1}\Delta_{1}(S)\cup\cdots\cup m_{\ell}\Delta_{1}(S)=\Delta_{1}(S)m_{1}^{-1}\cup\cdots\cup\Delta_{1}(S)m_{\ell}^{-1}$ (this will be proven below in Theorem 4.2). Since $\Delta_{1}(S)\subseteq D$ it follows that $G=m_{1}D\cup\cdots\cup m_{\ell}D=Dm_{1}^{-1}\cup\cdots\cup Dm_{\ell}^{-1}$.
hi $\square$
@@
The set $S$ in Theorem 4.7 has the property that $\Delta_{1}(S)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$ which means that for all $g$ such that $S\cap g^{-1}S\neq\emptyset$ it is true that $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g^{-1}S_{i}))>0$ for $1\leq i\leq k$. In fact one can show $S$ satisfies the stronger property that for all $g_{1},\dots,g_{n}\in G$ such that $S\cap(g_{1}^{-1}S)\cap\cdots\cap(g_{n}^{-1}S)\neq\emptyset$ it is true that $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g_{1}^{-1}S_{i})\cap\cdots\cap(g_{n}^{-1}S_{i}))>0$ for $1\leq i\leq k$. This result is reflected in the following theorem.
@@
Theorem 4.10. Let $G$ be a countably infinite amenable group with (left) Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$, where $\mathbf{G}$ is (left) tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is a set $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and if $g_{1},\ldots,g_{n}\in G$ are such that $S\cap(g_{1}^{-1}S)\cap\cdots\cap(g_{n}^{-1}S)\neq\emptyset$, then $\overline{d}_{\mathbf{F}_{i}}(S_{i}\cap(g_{1}^{-1}S_{i})\cap\cdots\cap(g_{n}^{-1}S_{i}))>0$ for $1\leq i\leq k$ [^10] In the special case where $k=1$, $G=\mathbb{Z}$, and the Følner sequences $\mathbf{G}$ and $\mathbf{F}_{1}$ are the standard Følner sequence $\{1,\ldots,N\}$, $N=1,2,3,...$, Theorem 4.2 reduces to a result of Ellis which was proven in [5, Corollary 2.1.1] and [19, Theorem 3.20]. It’s worth mentioning that both proofs of Ellis’ result rely on the pointwise ergodic theorem. Indeed, the proof of [5, Corollary 2.1.1] relies on Theorem 2 (see Remark 2) and the proof of [19, Theorem 3.20] relies on [19, Proposition 3.7] which establishes the existence of generic points for ergodic continuous maps of compact spaces with the help of the pointwise ergodic theorem.
@@
Utilizing the fact that every Følner sequence has a tempered Følner subsequence (see [22, Proposition 1.4]), it is straightforward to obtain the following variant of Theorem 4.2 with the slightly weaker assumption that the Følner sequence $\mathbf{G}$ may not be tempered.
@@
Theorem 4.11. Let $G$ be a countably infinite amenable group and let $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $G$. If $S_{1},\ldots,S_{k}\subseteq G$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there is $S\subseteq G$ such that $\overline{d}_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$, where
@@
To complete the logical circle, we still owe the reader the proof of Theorem 4.2 which was used in the proof of Theorem 4.2. Before proceeding, we need to introduce some preliminary definitions and prove some necessary lemmas. Let $G$ be a semigroup and let $S\subseteq G$. The set $S$ is left syndetic if there are $c_{1},\ldots,c_{k}\in G$ such that $c_{1}^{-1}S\cup\cdots\cup c_{k}^{-1}S=G$. Similarly $S$ is right syndetic if there are $d_{1},\dots,d_{r}\in G$ such that $Sd_{1}^{-1}\cup\dots\cup Sd_{r}^{-1}=G$. $S$ is right thick if for every finite $F\subseteq G$, there exists $g\in G$ for which $Fg\subseteq S$.
The next three lemmas are standard. We supply the proofs for the convenience of the reader.
@@
Proof. If $T=G$, then the result is clear. So suppose that $T\neq G$ and fix $t\in G\setminus T$. We inductively construct an injective sequence $s_{1},s_{2},...$ in $G$ with our desired property. Pick arbitrary $s_{1}\in G$. Now let $n\in\mathbb{N}$ be arbitrary and suppose that we have chosen pairwise distinct $s_{1},...,s_{n}\in G$ with $\{s_{i}^{-1}s_{j}:1\leq i<j\leq n\}\subseteq T$. Now $F=\left\{s_{1}^{-1},...,s_{n}^{-1},ts_{1}^{-1},...,ts_{n}^{-1}\right\}$ is finite, so take $s_{n+1}\in G$ such that $Fs_{n+1}\subseteq T$. Since $t\not\in T$, we have $s_{n+1}\neq s_{j}$ for any $1\leq j\leq n$. Furthermore one can now see that $\{s_{i}^{-1}s_{j}:1\leq i<j\leq n+1\}\subseteq T$, so by induction this gives us an infinite sequence $s_{1},s_{2},...$ with the desired properties. $\square$
@@
Proof. Let $F\subseteq G$ be finite and non-empty, say $F=\{g_{1},\ldots,g_{k}\}$. Then $g_{1}^{-1}S\cup\cdots\cup g_{k}^{-1}S\neq G$ by non-syndeticity of $S$, so pick $a\in G\setminus(g_{1}^{-1}S\cup\cdots\cup g_{k}^{-1}S)$. It follows that for all $g\in F$ we have $ga\notin S$, hence $Fa\subseteq G\setminus S=T$, which implies that $T$ is right thick as desired. $\square$
Lemma 4.14. Let $G$ be a countably infinite amenable group and let $\mathbf{F}$ be a Følner sequence in $G$. Let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. If $(s_{n})_{n=1}^{\infty}$ is an injective sequence in $G$, then there are $i<j$ such that $s_{i}^{-1}s_{j}\in\Delta_{1}(S)$.
Proof. Let $(s_{n})_{n=1}^{\infty}$ be an injective sequence in $G$. By the pigeonhole principle, there must be some $i<j$ for which $\overline{d}_{\mathbf{F}}((s_{i}S)\cap(s_{j}S))>0$. It follows that there are $u,v\in S$ such that $s_{j}u=s_{i}v$, thus $(s_{i}^{-1}s_{j})u=v$, and therefore $u\in S\cap(s_{i}^{-1}s_{j})^{-1}S$. It follows that $S\cap(s_{i}^{-1}s_{j})^{-1}S\neq\emptyset$, therefore $s_{i}^{-1}s_{j}\in\Delta_{1}(S)$. $\square$
We can now prove Theorem 4.2, thereby completing the proof of Theorem 4.2.
Theorem 4.15. Let $G$ be a countably infinite amenable group, let $\mathbf{F}$ be a Følner sequence in $G$, let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$, and let $E=\Delta_{1}(S)$. Then there exist $b_{1},...,b_{\ell}\in G$ with $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$ for which $b_{1}E\cup\cdots\cup b_{\ell}E=Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}=G$ (that is, $E$ is both left and right syndetic with the same “syndeticity constant” $\ell$).
Proof. By Lemma 4.2, for every injective sequence $(s_{n})_{n=1}^{\infty}$ in $G$ there are $i<j$ such that $s_{i}^{-1}s_{j}\in E$. Then by Lemma 4.2 it follows that $G\setminus E$ cannot be right thick, and so $E$ is left syndetic by Lemma 4.2, which implies that there are $c_{1},...,c_{k}\in G$ such that $c_{1}E\cup\cdots\cup c_{k}E=G$.
For any additional $g\in G$ with $g\notin\{c_{1},\ldots,c_{k}\}$ we have $g\in G=c_{1}E\cup\cdots\cup c_{k}E$, so $g\in c_{j}E$ for some $j\in\left\{1,...,k\right\}$. Hence $c_{j}^{-1}g\in E$, which means that $S\cap(c_{j}^{-1}g)^{-1}S\neq\emptyset$, so there are $u,v\in S$ such that $c_{j}^{-1}gu=v$. Thus $c_{j}\{u,v\}\cap g\{u,v\}\neq\emptyset$. It follows that there is a maximal subset $\left\{b_{1},...,b_{\ell}\right\}$ in $G$ such that, for all finite $F\subseteq S$, we have $b_{i}F\cap b_{j}F=\emptyset$ when $i\neq j$, and if $g\notin\{b_{1},\ldots,b_{\ell}\}$, then there is some finite $F\subseteq S$ such that $b_{i}F\cap gF\neq\emptyset$ for some $1\leq i\leq\ell$.
Write $\mathbf{F}=(F_{n})_{n=1}^{\infty}$ and for all $n\in\mathbb{N}$ put $S(n)=|A_{n}|$, where $A_{n}\vcentcolon=S\cap F_{n}$. By assumption, for all $n\in\mathbb{N}$, the sets $b_{1}A_{n},...,b_{\ell}A_{n}$ are disjoint. Hence,
@@
By taking $n\to\infty$ in formula (4.14) and using the fact that F is a Følner sequence, it follows that $\ell\leq 1/\overline{d}_{\mathbf{F}}(S)$.
@@
Lastly, note that if $d\in E$, then $d^{-1}\in E$. For all $g\in G$ we have $g^{-1}\in b_{1}E\cup\cdots\cup b_{\ell}E$, hence $b_{i}^{-1}g^{-1}=d$ for some $d\in E$ and so $gb_{i}=d^{-1}\in E$, thus $g\in Eb_{i}^{-1}$. This shows that $G=Eb_{1}^{-1}\cup\cdots\cup Eb_{\ell}^{-1}$. $\square$
4.3 Extending Theorem 4.2 to Cancellative Amenable Semigroups
In this subsection we will obtain a variant of Theorem 4.2 for cancellative amenable semigroups, namely Theorem 4.3. We will do so by utilizing the classical fact that cancellative amenable semigroups can be embedded into groups. Let $G$ be a cancellative amenable semigroup. It is known that there is a group $\widetilde{G}$ and an embedding $\varphi\colon G\to\widetilde{G}$ such that $\widetilde{G}=\left\{\varphi(s)\varphi(t)^{-1}:s,t\in G\right\}$ (see [15, Theorem 1.23] and [23, Proposition 1.23]). Moreover, the group $\widetilde{G}$ is unique (see [15, Theorem 1.25] for the precise formulation). With this in mind, we henceforth, when convenient, identify every element $g\in G$ with $\varphi(g)\in\widetilde{G}$ and call $\widetilde{G}$ the group of quotients of $G$.
@@
Theorem 4.16 ([3, Theorem 2.12]). Let $G$ be a countably infinite amenable cancellative semigroup. Then any Følner sequence $(F_{N})_{N=1}^{\infty}$ in $G$ is a Følner sequence in its group of quotients $\widetilde{G}$.
Remark 4.17. Let $G$ be a countably infinite amenable cancellative semigroup and let $\mathbf{F}=(F_{N})_{N=1}^{\infty}$ be a Følner sequence in $G$, and hence in $\widetilde{G}$. Let $S^{\prime}\subseteq\widetilde{G}$ and let $S\vcentcolon=S^{\prime}\cap G$. We have
@@
Moreover, if one of $d_{\mathbf{F}}(S)$ or $d_{\mathbf{F}}(S^{\prime})$ exists, then both exist and $d_{\mathbf{F}}(S)=d_{\mathbf{F}}(S^{\prime})$.